# Unified Stochastic Framework for Neural Network Quantization and Pruning

### v2

@@ 5. One-bit Quantization
We are now ready to first describe the selection of $\mathcal{T}$ and $\theta$ for one-bit quantization and then provide the corresponding error analysis for running Algorithm 1 on a single layer of a neural network. Notably, our analysis extends with minimal additional effort to finite multi-bit quantization, a consideration we leave to the interested reader.
5.1. Choice of $\mathcal{T}$ and $\theta$ for one-bit quantization
Suppose we want to quantize an $L$-layer perceptron which has all its weights bounded in absolute value by $K>0$. We use $\mathcal{A}=\{-2K,2K\}$ as the alphabet. Then, we can choose $\mathcal{T}$ to be the stochastic scalar quantizer $\mathcal{Q}:\mathbb{R}\rightarrow\mathcal{A}$, with
$\mathcal{Q}(z)=\begin{cases}2K&\text{with probability $p_{z}$ if $|z|\leq 2K$}\\ -2K&\text{with probability $1-p_{z}$ if $|z|\leq 2K$}\\ \text{sgn}(z)\cdot 2K&\text{if $|z|>2K$},\end{cases}$
where $p_{z}=\tfrac{1}{2}+\tfrac{z}{4K}$. Moreover, we choose the threshold $\theta=K$. It is straightforward to verify that the operator $\mathcal{Q}$ satisfies the assumptions in Theorem 4.2 with $M=4K$.
@@ 5.2. Error analysis
Let $C\geq 1$. Suppose that $Q$ is the quantized weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic quantizer $\mathcal{Q}$. Then, the resulting error satisfies
$\max_{i,j}|\rho(XW)_{ij}-\rho(XQ)_{ij}|\leq 4K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|$
@@
Let $\kappa:=4K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|$. Consider any neuron (column) $w\in\mathbb{R}^{\mathbb{N}_{0}}$ from $W$. Let $q$ and $\widetilde{u}_{i}$ be as defined in Subsection 4.1. We run Algorithm 2 independently for each $w$. We can deduce
[…]
$=\mathbb{P}\big{(}\text{$|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$ for all $t$ and }\|\widetilde{u}_{N}\|_{\infty}\leq\kappa\big{)}$
$=1-\mathbb{P}(\text{$|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|>K$ for some $t$ or }\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa)$
$\geq 1-\sum_{t}\mathbb{P}\Big{(}\Big{|}\dfrac{\langle\widetilde{u}_{t},X_{t+1}\rangle}{C\|X_{t+1}\|^{2}}\Big{|}>K\Big{)}-\mathbb{P}(\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa).$
By Theorem 4.2 and Lemma 4.3, b, we have
$\dfrac{\langle\widetilde{u}_{t},X_{t+1}\rangle}{C\|X_{t+1}\|^{2}}\prec_{cx}\mathcal{N}\Big{(}0,\dfrac{\beta_{t}}{C^{2}\|X_{t+1}\|^{2}}\Big{)},$
where $\beta_{t}=8C\pi K^{2}\max_{1\leq i\leq t}\|X_{i}\|^{2}.$ Applying Lemma 4.3, g, we obtain
$\mathbb{P}\Big{(}\Big{|}\dfrac{\langle\widetilde{u}_{t},X_{t+1}\rangle}{C\|X_{t+1}\|^{2}}\Big{|}>K\Big{)}\leq\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t+1}\|^{2}}{32\pi\max_{1\leq i\leq t}\|X_{i}\|^{2}}}\Big{)}.$ (1)
For $\mathbb{P}(\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa)$, we use the second part of Lemma 4.3, g with $\gamma=\sqrt{2}mN_{0}^{-p}$ to get
$\mathbb{P}(\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa)\leq\sqrt{2}mN_{0}^{-p}.$ (2)
Thus using (1) and (2), we obtain
[…]
$\geq 1-\sum_{t=2}^{N_{0}}\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t}\|^{2}}{32\pi\max_{1\leq i\leq t-1}\|X_{i}\|^{2}}}\Big{)}-\sqrt{2}mN_{0}^{-p}.$
@@
$\mathbb{P}(\|\rho(Xw)-\rho(Xq)\|_{\infty}\leq\kappa)$ (3)
@@
by the definition of $\widetilde{u}_{N_{0}}$ and the the Lipschitz property of the activation function $\rho$. Inequality (3) holds true for each column of $W\in\mathbb{R}^{m\times N_{1}}$. Taking a union bound on the probability completes the proof. ∎
@@
This result demonstrates that by choosing $C$ proportional to $\log(N_{0}N_{1})$ the error associated with quantization scales as ${\operatorname{polylog}N_{0}}$, with high probability. This stands in sharp contrast to the error from naive rounding algorithms, where each entry of $w$ is rounded to the nearest element of $\mathcal{A}$. In the naive case, the error scales as $\sqrt{N_{0}}$.
@@ 6. Pruning
In this section, we specify the choice of $\mathcal{T}$ and the threshold $\theta$ used for pruning, and we present the error analysis of applying Algorithm 1 to a single layer of a neural network.
6.1. Choice of $\mathcal{T}$ and $\theta$ for Pruning
@@
Here $\xi_{z}$ is a random variable that is set to 0 with probability $1-\tfrac{1}{(c+1/2)K}|z|$ and is sampled from $\operatorname{sgn}(z)\cdot{U}[cK,K]$ with probability $\tfrac{1}{(c+1/2)K}|z|$. Additionally, we set $\theta$ to be $\infty$. It is straightforward to verify that the operator $\mathcal{S}$ satisfies the assumptions in Theorem 4.2 with $M$ being $K$.
@@ 6.2. Error analysis
Let $C\geq 1$. Suppose that $Q$ is the pruned weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic pruning operator $\mathcal{S}$. Then, the resulting error satisfies
@@
Consider any neuron (column) $w\in\mathbb{R}^{\mathbb{N}_{0}}$ from $W$. Let $q$ and $\widetilde{u}_{i}$ be defined as in Subsection 4.1. We run Algorithm 2 independently for each $w$. In this case, $\widetilde{u}_{i}=u_{i}$ for all $i$.
@@
$u_{j}\prec_{cx}\mathcal{N}(0,\beta_{j}I)\text{ for $1\leq j\leq N_{0}$},$
where $\beta_{j}=\tfrac{C\pi K^{2}}{2}\max_{1\leq i\leq j}\|X_{i}\|^{2}.$
Applying Lemma 4.3, g with $\gamma=\sqrt{2}mN_{0}^{-p}$, we obtain
@@
The inequality (5) holds for any column $w$ of $W\in\mathbb{R}^{N_{0}\times N_{1}}$. Applying a union bound on the probability completes the proof. ∎
@@ 7. Quantization with pruning
In this section, we show that our methods can handle combinations of operators, and provide the choices for $\mathcal{T}$ and the threshold $\theta$ used in quantization with pruning, followed by an error analysis for applying Algorithm 1 to a single layer of a neural network.
7.1. Choice of $\mathcal{T}$ and $\theta$ for Quantization with Pruning
Consider an $L$-layer perceptron with all weights bounded by $K>0$ in absolute value. Given $c>0$, and alphabet $\mathcal{A}=\{-2K,0,2K\}$, our goal is to encourage weights below $cK$ to be set to 00 and then quantize the weights using $\mathcal{A}$. Let $\mathcal{S}$ be the operator defined in Section 6. Also, we define the following stochastic rounding operator $\mathcal{Q}:\mathbb{R}\rightarrow\mathcal{A}$
$\mathcal{Q}(z)=\begin{cases}\lfloor\tfrac{z}{2K}\rfloor\cdot 2K&\text{with probability $p_{z}$ if $|z|\leq 2K$}\\ \left(\lfloor\tfrac{z}{2K}\rfloor+1\right)\cdot 2K&\text{with probability $1-p_{z}$ if $|z|\leq 2K$}\\ \text{sgn}(z)\cdot 2K&\text{if $z>2K$},\end{cases}$
where $p_{z}=1-\tfrac{z}{2K}+\lfloor\tfrac{z}{2K}\rfloor$. We then define $\mathcal{T}=\mathcal{Q}\circ\mathcal{S}$ with the threshold $\theta=K$. As before, it is straightforward to verify that the operator $\mathcal{S}$ satisfies the assumptions of Theorem 4.2 with $M=2K$.
@@ 7.2. Error analysis
Let $C\geq 1$. Suppose that $Q$ is the pruned and quantized weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic operator $\mathcal{Q}\circ\mathcal{S}$. Then, the resulting error satisfies
$\max_{i,j}|\rho(XW)_{ij}-\rho(XQ)_{ij}|\leq 2K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|$
@@
The proof of this Proposition is essentially identical to that of Proposition 1, so we omit the details.

### v3

@@ 5. One-bit Quantization
We are now ready to first describe the selection of $\mathcal{T}$ for one-bit quantization and then provide the corresponding error analysis for running Algorithm 1 on a single layer of a neural network. Notably, our analysis extends with minimal additional effort to finite multi-bit quantization, a consideration we leave to the interested reader.
5.1. Choice of $\mathcal{T}$ for one-bit quantization
Suppose we want to quantize an $L$-layer perceptron which has all its weights strictly bounded in absolute value by $K>0$, and our goal is to achieve one-bit quantization. We begin by considering an infinite alphabet $\mathcal{A}=\{\dots,-10K,-6K,-2K,2K,6K,10K,\dots\}$. Then, we choose the operator $\mathcal{T}$ to be the stochastic scalar quantizer $\mathcal{Q}:\mathbb{R}\rightarrow\mathcal{A}$, with
$\mathcal{Q}(z)=\begin{cases}\left(\lfloor\frac{z}{4K}\rfloor\right)\cdot 4K&\text{with probability $p_{z}$}\\ \left(\lceil\frac{z}{4K}\rceil\right)\cdot 4K&\text{with probability $1-p_{z}$},\\ \end{cases}$
where $p_{z}=\lceil\frac{z}{2K}\rceil-\frac{z}{2K}$.
It is straightforward to verify that the operator $\mathcal{Q}$ satisfies the assumptions in Theorem 4.2 with $M=4K$. We will show that with this choice of operator, with high probability, Algorithm 1 gives logarithmic reconstruction error with respect to the input dimension, and, crucially, only ever uses two values from the alphabet: $-2K$ and $2K$. That is, although the quantizer is defined over an infinite alphabet, the output of the algorithm effectively requires only 1-bit quantization with high probability, while maintaining small errors.
@@ 5.2. Error analysis
Let $C\geq 1$ and fix any $p\geq 1$. Suppose that $Q$ is the quantized weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic quantizer $\mathcal{Q}$. Then, the resulting error satisfies
$\max_{i,j}|\rho(XW)_{ij}-\rho(XQ)_{ij}|\leq 4K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N_{0}}\|X_{i}\|$
and
$\max_{i,j}|Q_{i,j}|\leq 2K$
@@
Let $\kappa:=4K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|$. Consider any neuron (column) $w\in\mathbb{R}^{N_{0}}$ from $W$. Let the quantized neuron $q$ and accumulated error $u_{t}$ be as defined in Subsection 4.1. By Theorem 4.2, we have $u_{t}\prec_{cx}\mathcal{N}(0,\beta_{t}I)$, where $\beta_{t}$ is defined in Lemma 4.5. Recall that at each step in Algorithm 1, the quantized weight is $q_{t}=\mathcal{Q}\left(w_{t}+\tfrac{\langle u_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}\right).$ Let $B_{t}$ be the event $\{|q_{t}|>2K\}$. We know $B_{t}\subseteq A_{t}=\left\{|\tfrac{\langle u_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|>K\right\}$ since $|w_{t}|<K$. To obtain the result in the statement, it suffices to bound the following failure probability
($\mathbb{P}\left(\left\{\|u_{N_{0}}\|_{\infty}>\kappa\right\}\cup A_{1}\cup\dots\cup A_{N_{0}}\right).$
We now bound the probability of each set individually. First, $\mathbb{P}(A_{1})$ is 0 since $u_{0}=0$. We then control $\mathbb{P}(A_{t})$ for $t=2,\dots,N_{0}$. By Theorem 4.2 and Lemma 4.3, b, we have
$\dfrac{\langle u_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}\prec_{cx}\mathcal{N}\Big{(}0,\dfrac{\beta_{t-1}}{C^{2}\|X_{t}\|^{2}}\Big{)},$
where $\beta_{t-1}=8C\pi K^{2}\max_{1\leq i\leq t-1}\|X_{i}\|^{2}.$ Applying Lemma 4.3, f, we obtain
$\mathbb{P}\Big{(}\Big{|}\dfrac{\langle u_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}\Big{|}>K\Big{)}\leq\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t}\|^{2}}{32\pi\max_{1\leq i\leq t-1}\|X_{i}\|^{2}}}\Big{)}.$ (1)
For ($\mathbb{P}\left(\left\{\|u_{N_{0}}\|_{\infty}\geq\kappa\right\}\right)$, we use the second part of Lemma 4.3, f with $\gamma=\sqrt{2}mN_{0}^{-p}$ to get
$\mathbb{P}(\|u_{N_{0}}\|_{\infty}>\kappa)\leq\sqrt{2}mN_{0}^{-p}.$ (2)
Then we use apply a union bound on the failure probability to obtain an upper bound
($\mathbb{P}\left(\left\{\|u_{N_{0}}\|_{\infty}>\kappa\right\}\cup A_{1}\cup\dots\cup A_{N_{0}}\right)\leq\sqrt{2}mN_{0}^{-p}+\sum_{t=2}^{N_{0}}\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t}\|^{2}}{32\pi\max_{1\leq i\leq t-1}\|X_{i}\|^{2}}}\Big{)}.$
@@
$\mathbb{P}(\|\rho(Xw)-\rho(Xq)\|_{\infty}\leq\kappa\text{ and }\|q\|_{\infty}\leq 2K)$ (3)
@@
by the definition of $u_{N_{0}}$ and the the Lipschitz property of the activation function $\rho$. Inequality (3) holds true for each column of $W\in\mathbb{R}^{m\times N_{1}}$. Taking a union bound over all the columns completes the proof.
@@
This proposition ensures that each entry of $|Q|$ is bounded by $2K$ with high probability. Moreover, since $Q$ takes values in the discrete alphabet $\mathcal{A}$ and the only elements of $\mathcal{A}$ within $[-2K,2K]$ are $-2K$ and $2K$, this implies that each entry of $Q$ must be either $-2K$ or $2K$. Hence, although the quantizer is defined over a larger alphabet, the quantized weights are effectively supported on just two values, realizing 1-bit quantization in practice.
@@
This result demonstrates that, for one neuron $w$, by choosing $C$ proportional to $\log(N_{0})$, the Euclidean norm error $\|Xw-Xq\|$ associated with quantization scales as $K\sqrt{m}\cdot{\operatorname{polylog}N_{0}}\cdot\max_{i}\|X_{i}\|$ with high probability, where $m$ is the number of data points. In the case when $m>N$, one can modify the argument in the proof of Proposition 1 via an SVD of the data matrix $X=U\Sigma V^{\top}$. To be more specific, one can prove that running Algorithm 2 with $X$ is equivalent to running it with $\Sigma V^{\top}$. Moreover, $\|X(w-q)\|=\|\Sigma V^{\top}(w-q)\|$, thereby yielding an error of order $K\sqrt{r}\cdot{\operatorname{polylog}N_{0}}\cdot\max_{i}\|X_{i}\|$, where $r\leq N_{0}$ is the rank of $X$. In summary, the error scales like $K\sqrt{\min\{m,N_{0}\}}\cdot{\operatorname{polylog}N_{0}}\cdot\max_{i}\|X_{i}\|$. This stands in sharp contrast to the Euclidean norm error from the naive round-to-nearest (RTN) algorithm, where each entry of $w$ is rounded to the nearest element of $\mathcal{A}$. In that case, the only error bound, in general, is via triangle inequalities and leads to a Euclidean norm error $\|X(w-q)\|\leq\sum_{i=1}^{N_{0}}\|X_{i}\||w_{i}-q_{i}|\leq 2K\cdot N_{0}\cdot\max_{i}\|X_{i}\|$.
@@ 6. Pruning
In this section, we specify the choice of $\mathcal{T}$ used for pruning, and we present the error analysis of applying Algorithm 1 to a single layer of a neural network.
6.1. Choice of $\mathcal{T}$ for Pruning
@@
Here $\xi_{z}$ is a random variable that is set to 0 with probability $1-\tfrac{2}{(c+1)K}|z|$ and is sampled from $\operatorname{sgn}(z)\cdot{U}[cK,K]$ with probability $\tfrac{2}{(c+1)K}|z|$. It is straightforward to verify that the operator $\mathcal{S}$ satisfies the assumptions in Theorem 4.2 with $M$ being $K$.
@@ 6.2. Error analysis
Let $C\geq 1$ and fix any $p\geq 1$. Suppose that $Q$ is the pruned weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic pruning operator $\mathcal{S}$. Then, the resulting error satisfies
@@
Consider any neuron (column) $w\in\mathbb{R}^{\mathbb{N}_{0}}$ from $W$. Let $q$ and $u_{i}$ be defined as in Subsection 4.1. We run Algorithm 2 independently for each $w$.
@@
$u_{t}\prec_{cx}\mathcal{N}(0,\beta_{t}I)\text{ for $1\leq t\leq N_{0}$},$
where $\beta_{t}=\tfrac{C\pi K^{2}}{2}\max_{1\leq i\leq t}\|X_{i}\|^{2}.$
Applying Lemma 4.3, f with $\gamma=\sqrt{2}mN_{0}^{-p}$, we obtain
@@
The inequality (5) holds for any column $w$ of $W\in\mathbb{R}^{N_{0}\times N_{1}}$. Applying a union bound over all the columns completes the proof. ∎
@@ 7. Quantization with pruning
In this section, we show that our methods can handle combinations of operators, and provide the choices for $\mathcal{T}$ used in joint quantization with pruning, followed by an error analysis for applying Algorithm 1 to a single layer of a neural network.
7.1. Choice of $\mathcal{T}$ for Quantization with Pruning
Consider an $L$-layer perceptron with all weights strictly bounded by $K>0$ in absolute value. We again begin with an infinite alphabet $\mathcal{A}=\{\dots,-6K,-4K,-2K,0,2K,4K,6K,\dots\}$. Let $\mathcal{S}$ be the pruning operator defined in (4), with $c>0$, and let $\mathcal{Q}:\mathbb{R}\rightarrow\mathcal{A}$ be the stochastic scalar quantizer with
$\mathcal{Q}(z)=\begin{cases}\left(\lfloor\frac{z}{2K}\rfloor\right)\cdot 2K&\text{with probability $p_{z}$}\\ \left(\lceil\frac{z}{2K}\rceil\right)\cdot 2K&\text{with probability $1-p_{z}$},\\ \end{cases}$
where $p_{z}=\lceil\frac{z}{2K}\rceil-\frac{z}{2K}$. Our goal is to encourage weights below $cK$ to be set to 00 and then quantize the weights using $\mathcal{A}$. To that end, we define $\mathcal{T}=\mathcal{Q}\circ\mathcal{S}$.
@@ 7.2. Error analysis
Let $C\geq 1$, and fix any $p\geq 1$. Suppose that $Q$ is the pruned and quantized weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic operator $\mathcal{T}=\mathcal{Q}\circ\mathcal{S}$. Then, the resulting error satisfies
$\max_{i,j}|\rho(XW)_{ij}-\rho(XQ)_{ij}|\leq 2K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N_{0}}\|X_{i}\|$
and
$\max_{i,j}|Q|_{ij}\leq 2K$
@@
Proof.
First, we verify that the operator $\mathcal{T}$ satisfies the assumptions of Theorem 4.2 with $M=2K$.
We begin by showing that for any $z\in\mathbb{R}$, the error $|\mathcal{T}(z)-z|$ is bounded by $2K$. Indeed, if $|z|>cK$, then $\mathcal{S}(z)=z$, so $\mathcal{T}(z)=\mathcal{Q}(z)$, which rounds $z$ up or down to an adjacent alphabet element; since the grid size is $2K$, the rounding error is at most $2K$. If $|z|\leq cK$, then $\mathcal{S}(z)$ maps $z$ to a value in $[-K,K]$ with the same sign. Without loss of generality, suppose $z\in[0,K]$, so $\mathcal{S}(z)\in[0,K]$ and $\mathcal{Q}(\mathcal{S}(z))\in[0,2K]$. Hence, both $z$ and $\mathcal{T}(z)$ lie in $[0,2K]$, and the total deviation remains at most $2K$.
For the expectation, we consider two independent probability spaces $(\Omega_{1},\mathcal{E}_{1},\mathbb{P}_{1})$ and $(\Omega_{2},\mathcal{E}_{2},\mathbb{P}_{2})$, which govern the randomness of the operators $\mathcal{Q}_{\omega_{1}}$ and $\mathcal{S}_{\omega_{2}}$ respectively, with $\omega_{1}\in\Omega_{1}$ and $\omega_{2}\in\Omega_{2}$. Jointly, the probability space for $\mathcal{T}$ is the direct product of these two spaces. Then, for any $z\in\mathbb{R}$,
[…]
$=\mathbb{E}_{\omega_{2}}[\mathcal{S}_{\omega_{2}}(z)]=z.$
The third equality follows from the unbiasedness of $\mathcal{Q}$, and the final equality from the unbiasedness of $\mathcal{S}$. The rest of the proof proceeds identically to that of Proposition 1, so we omit the details. ∎